# “Always Nice and Confident, Sometimes wrong”: Developer’s Experiences Engaging Generative AI Chatbots Versus Human-Powered Q&A Platforms

### v2

@@ Abstract.
Software engineers have historically relied on human-powered Q&A platforms, like Stack Overflow (SO), as coding aids. With the rise of generative AI, developers have adopted AI chatbots, such as ChatGPT, in their software development process. Recognizing the potential parallels between human-powered Q&A platforms and AI-powered question-based chatbots, we investigate and compare how developers integrate this assistance into their real-world coding experiences by conducting thematic analysis of Reddit posts. Through a comparative study of SO and ChatGPT, we identified each platform’s strengths, use cases, and barriers. Our findings suggest that ChatGPT offers fast, clear, comprehensive responses and fosters a more respectful environment than SO. However, concerns about ChatGPT’s reliability stem from its overly confident tone and the absence of validation mechanisms like SO’s voting system. Based on these findings, we synthesized the design implications for future GenAI code assistants and recommend a workflow leveraging each platform’s unique features to improve developer experiences.
Q&A platform, Programming assistance, Generative AI
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXccs: Human-centered computing User studiesccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Natural language interfaces
@@ 2.1. Q&A platform for programming practices
In the past few decades, social media has dramatically changed the landscape of software engineering, with Q&A platforms as one of the most recent tools emerging around 2010’s and widely adapted by programmers (Storey et al., 2014). Having improved response time significantly compared to traditional communication methods, social media is attractive for software developers (Mamykina et al., 2011b). The first prototype of Q&A sites was established by Ackerman’s Answer Garden (Ackerman and Malone, 1990), and gradually evolved and exposed to the general population. In 2005, both Reddit and Yahoo! Answers, 2 popular Q&A platforms were released to the public, with Reddit still being one of the most popular social media platform as of 2023. Software engineers quickly started to make use of these platforms, established Stack Exchange Network which then transformed to Stack Overflow, a Q&A websites that focused on programming related questions. Research has followed to learn more about this transition and revolution. Storey et al. well described the revolution of social media in software engineering from non-digital, digital to socially enabled (including Q&A platform) (Storey et al., 2014). Squire et al. conducted the first study to assess the quality of developer support provided by Q&A sites such as SO, in comparison with previous tools used for developer support, such as mailing list (Squire, 2015). According to Mamykina et al., some software developers started to believe that Q&A platforms such as SO had replaced web search/forums as their main source of finding answers to their programming problems (Mamykina et al., 2011a). With this trend happening, we are curious about how Q&A platforms will continue to evolve to better fit programmers’ needs. Hence we chose SO, one of the most popular Q&A site that concentrating on programming related question, as our proxy for this research.
@@ 2.1.1. Stack Overflow
Created in 2008, SO quickly became the most widely used Q&A platform for software engineers in the world with about 50 million visits monthly (Nasehi et al., 2012; SEd, 2023). SO is centered around nine design decisions: voting, tags, editing, badges, karma, pre-search, search engine optimization, user interface, and critical mass (Treude et al., 2011). As such a well-developed and popular platform, SO triggered numerous research to conduct study on this platform. Mamykina1 et al. examined what contributed to SO’s success, specifically engaged with the founder of this platform (Mamykina et al., 2011a). As a question-initiated system, much research focuses on analyzing the questions asked on SO, with Treude et al.’s paper in 2011 being one of the mostly cited work around SO (Treude et al., 2011). They summarized 7 categories of questions: how-to, discrepancy, environment, error, decision help, conceptual, and review. In addition to the high-level summary, researchers also summarized the detailed use cases of SO (ins, 2023; Antelmi et al., 2023), discussed ’how to ask the right questions’ (Ahmad et al., 2018; Yang et al., 2014), as well as analyzing the characteristics of the answerslike response time (Mamykina et al., 2011a) and attributes of recognized answers (Nasehi et al., 2012). Other aspects of SO like reward system(Wang et al., 2018), novice experience (Chatterjee et al., 2020), and social norms (Cheriyan et al., 2021), were also discussed in various studies. These studies shed light on different aspects of SO’s functioning and user behavior. In addition to analyzing data and deriving insights, researchers have also developed tools to automate various aspects of this process, including classifier (Beyer et al., 2018), datasets (Baltes et al., 2018), and code search and recommendation tool (Zagalsky et al., 2012). These tools contribute to the efficiency and comprehensiveness of SO research efforts.
@@ 2.2. Generative AI for coding
Transitioning from human-powered Q&A platforms, there is a growing trend of people increasingly embracing AI-powered tools to aid in their programming processes. According to SO’s sentiment report in 2023, 70% of the developers are already using or plan to use AI tools in their development process (sen, 2023). This trend has been significantly influenced by the advancements in LLM, particularly publicly available tools. One noteworthy example is Github Copilot, which was announced on June 29, 2021. Copilot, usually considered as an LLM-based code generation tool (Vaithilingam et al., 2022), offers various assistive features for programmers, including the conversion of code comments into executable code and autocompletion for code segments, repetitive sections, entire methods, and functions (cop, 2023).
Although it is a relatively new topic, researchers have already begun exploring programmers’ experiences with LLM-powered code generation tools. For instance, Vaithilingam et al. conducted a within-subjects user study with 24 participants to understand how programmers use and perceive Copilot (Vaithilingam et al., 2022). While these code generation tools offer helpful assistance, they cannot fulfill all the needs typically met by traditional question-based solutions. Alongside these tools, there has also been a rising adoption of Gen AI-powered chatbots for coding, a method that closely parallels traditional Q&A platforms. For instance, Rose et al. created the Programmer’s Assistant, a conversational prototype system, to explore conversational interactions rooted in code. Beyond small-scale research endeavors, the advent of ChatGPT spurred a larger trend of embracing AI chatbots for coding purposes. Launched on November 30, 2022, ChatGPT has made a substantial social impact. Users quickly began sharing their experiences and discussions on social media platforms, providing researchers with an opportunity to investigate user experiences.
Given the novelty of this topic, researchers have just begun planning studies, but there were already many interesting works that we could learn from. One primary area of interest is testing the accuracy of ChatGPT-generated answers. Kabir et al. conducted a study using SO posts as input and assessed ChatGPT’s accuracy in generating answers and found a 52% inaccuracies (Kabir et al., 2023). Despite inaccuracies, researchers still found users’ expressing willingness in using these GenAI tools (Lau and Guo, 2023).
@@
Beyond these tensions between platforms, recent research has also begun to question ChatGPT’s effectiveness in aiding coding tasks by comparing it with other methods. For example, Choudhuri et al. conducted an empirical study and discovered that ChatGPT did not enhance participants’ productivity or self-efficacy compared to traditional non-GenAI online resources (Choudhuri et al., 2024). Instead, it significantly heightened their frustration levels. While their findings highlight significant issues with GenAI coding assistance, these results are constrained by the small sample size of their lab-based experiment. We hold a more optimistic perspective and aim to observe long-term adoption of these tools in real-world settings to integrate the strengths of both methodologies. A similar view stemmed from Cheng et al., where they discussed how online communities shape developers’ trust in AI tools and explored ways to leverage community features to foster appropriate user trust in AI (Cheng et al., 2022). Rather than creating opposition, our aim is to explore opportunities for collaboration and improvement based on the unique affordances of each tool. To facilitate this discussion, we will compare ChatGPT with SO and propose strategies and solutions. We anticipate that analyzing Reddit posts will uncover long-term solutions adopted by programmers in real-world settings, providing valuable insights into how we might advance our approach.
@@ 3.2. Data filtering
During the phase of data exploration, we observed an overlap in the posts on the two identified themes. For example, many posts in r/ChatgptCoding also discussed SO. Recognizing this pattern, we combined all the post data from different channels together as our final dataset. We then proceeded with an initial data filtering such as eliminating duplicate posts. Subsequently, to enhance the precision of our dataset, manual filtering of irrelevant content was undertaken. This process involved researchers reviewing post content (selftext) and titles to ensure alignment. Ultimately, we reduced our dataset from over 4000 posts to 1,758 posts. The datasets are publicly available  [^1] https://figshare.com/s/74e3b2e45a352c864b15.
@@ 4.2. Analysis
We generated a wordcloud [^2] https://github.com/amueller/word_cloud to visually represent the narratives across all the posts [Fig. 2]. We also performed sentiment analysis on the posts using one of the state-of-the-art unsupervised sentiment prediction tool VADER [^3] https://github.com/cjhutto/vaderSentiment5 (Hutto and Gilbert, 2014). The sentiment analysis of the posts reveals a moderately positive tone throughout the dataset, with an average compound score of 0.22. Among these posts, 37.91% are categorized as positive, 53.70% as neutral, and 10.13% as negative. This distribution suggests that while the majority of the discussions are neutral or mildly positive, a significant portion still reflects varied sentiments, underlining the complexity of user experiences.
@@ 5. Qualitative results
In order to answer RQ1: In what ways are individuals employing GenAI chatbots to enhance their coding experience?, we would like to discuss three themes: Strengths, Use Cases and Barriers. The comprehensive codebooks are available in the tables present in Appendix.
@@ 5.1. Strengths
Many individuals have reported highly successful experiences using ChatGPT to enhance their coding endeavors. In terms of the overall user experience, ChatGPT’s adoption of natural language interactions contributes to an exceptionally barrier to entry. One user expressed this by stating, ”You will literally just describe the tool you need and you will get it.” Notably, even individuals who had previously given up on their programming pursuits expressed optimism about ChatGPT, believing it would help them ”see it through this time.” This gentle learning curve underscores ChatGPT’s potential in learning scenarios, particularly for novice users.
@@ 5.3. Barriers
One of the most prominent issues with ChatGPT’s answers is related to reliability. Specifically in coding scenarios, there have been instances where ChatGPT generated different fictitious codes each time it was asked the exact same question, undermining users’ trust in the system. To the surprise of many programmers, ChatGPT often fabricated non-existent libraries, commands, or citations. These red herrings, or false leads, can be particularly detrimental as it places a significant burden on users to verify each piece of information, potentially eroding trust in the entire solution. Users have also noted that even when ChatGPT provides incorrect or ambiguous answers, it does so with a ”calm and assured confidence,” highlighting a disparity between ChatGPT’s responses and users’ expectations, especially when users begin to treat it as a real assistant and apply human-like judgment. This unreliability is exacerbated by the fact that ChatGPT operates as a standalone program. As one user put it, ”ChatGPT only helps if you can judge whether its answer is correct, unlike an answer with 50 upvotes on SO, which you _know_ is correct.”
@@
Unsurprisingly, users have encountered limitations in terms of resource access, including constrained training sets, the absence of live resources after 2021, the inability to run ChatGPT locally, and the lack of integrated development environment (IDE) plugins. Additionally, there have been instances where ChatGPT refused to generate code, especially when specific trigger prompts were used. Users have also expressed the need for an ’adult’ version of ChatGPT with fewer restrictions, as many of the current limitations are related to ethical concerns.
As ChatGPT gained popularity, some users encountered difficulties accessing the system. These constraints included rate limitations, regional restrictions, financial implications, downtime, and system overload. Additionally, some users expressed concerns about copyright issues related to using ChatGPT’s generated code, particularly in professional settings.
@@
We noted that many Reddit posts were based on personal experiences, but one consensus emerged: ChatGPT was perceived as significantly more user-friendly than SO. SO has a long reputation of being toxic. ”The fact is, SO has a serious problem with bullying, one that SO itself has been trying really hard to mitigate, without success so far.”, from one Reddit post. In contrast, when using Chatgpt, users did not need to worry about asking ”stupid” or repeated questions.
@@
Use Cases Personal assistant + Learning potencials 7 categories (Treude et al., 2011)
@@ 6.1. Long-Term adoption of GenAI in coding: Reddit posts versus lab experiments
Our results complement experimental studies of coding experiences with ChatGPT and Copilot (Choudhuri et al., 2024; Vaithilingam et al., 2022). Prior work has also noted that Gen AI assistants excel at providing a starting point, are easy for beginners to use, and save programmers’ time. Perhaps the most interesting findings, however, are around differing results and conclusions. Choudhuri et al. reported negative outcomes from using ChatGPT for coding assistance, noting that it neither enhanced productivity nor self-efficacy and instead increased frustration (Choudhuri et al., 2024). However, in our analysis of Reddit posts, sentiment analysis revealed moderate positivity and the qualitative feedback from users presented a more mixed picture with both strengths and barriers. In contrast, long-term users might have had more time to adapt and find effective ways to utilize the tool, developing specific prompting strategies like “explain certain code examples as if it were explaining to a 10 year old” mentioned in one of the posts. These prior studies focused on task completion, which, while important, does not capture the long-term sentiments about these tools, which we were able to identify in the posts. In our findings, individuals who have used ChatGPT for coding assistance in real-world scenarios greatly appreciated the respectful and patient environment it fostered, contrasting with their experiences on platforms like SO. Access and copyright issues, often overlooked in lab studies, are incredibly crucial and significantly impact long-term usage, especially in professional settings.
Programmers have more time to explore different uses of the tools in-the-wild, not restricted to task completion as dictated by specific study designs. Many Reddit posts indicate that programmers use ChatGPT for coding-adjacent activities, such as generating documentation, unit tests, and managing edge cases. Other use cases, such as learning, could only be captured in long-term usage with ChatGPT, where users viewed it as a coding assistant and pair programming partner, and even used it to generate quizzes. Similarly, other emergent use cases, such as employing it as a code editor or for simulating operational systems, can be detected using our methodology.

### v3

# ”Always Nice and Confident, Sometimes Wrong”: Developer’s Experiences Engaging Large Language Models (LLMs) Versus Human-Powered Q&A Platforms for Coding Support

@@ Abstract.
Software engineers have historically relied on human-powered Q&A platforms like Stack Overflow (SO) as coding aids. With the rise of generative AI, developers have started to adopt AI chatbots, such as ChatGPT, in their software development process. Recognizing the potential parallels between human-powered Q&A platforms and AI-powered question-based chatbots, we investigate and compare how developers integrate this assistance into their real-world coding experiences by conducting a thematic analysis of 1700+ Reddit posts. Through a comparative study of SO and ChatGPT, we identified each platform’s strengths, use cases, and barriers. Our findings suggest that ChatGPT offers fast, clear, comprehensive responses and fosters a more respectful environment than SO. However, concerns about ChatGPT’s reliability stem from its overly confident tone and the absence of validation mechanisms like SO’s voting system. Based on these findings, we synthesized the design implications for future GenAI code assistants and recommend a workflow leveraging each platform’s unique features to improve developer experiences.
Developer Support, Programming Assistance, Generative AI, Large Language Model, Human-AI Collaboration, Q&A Platform, Stack Overflow, ChatGPT, User Experience, Social Media, Reddit, Data Mining, Thematic Analysis
copyright: acmcopyrightjournalyear: 2025ccs: Software and its engineering Collaboration in software developmentccs: Human-centered computing HCI theory, concepts and modelsccs: Human-centered computing Collaborative and social computing theory, concepts and paradigmsccs: Social and professional topicsccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Development frameworks and environmentsccs: Human-centered computing User studiesccs: Human-centered computing Empirical studies in collaborative and social computingccs: Human-centered computing Natural language interfacesccs: Human-centered computing Collaborative interaction
@@ 2.1. Q&A Platform for Programming Practices
In the past few decades, social media has dramatically changed the landscape of software engineering, with Q&A platforms as one of the most recent tools emerging around 2010’s and widely adapted by programmers (Storey et al., 2014). Having improved response time significantly compared to traditional communication methods, social media is attractive for software developers (Mamykina et al., 2011b). The first prototype of Q&A sites was established by Ackerman’s Answer Garden (Ackerman and Malone, 1990), and gradually evolved and exposed to the general population. In 2005, both Reddit and Yahoo! Answers, 2 popular Q&A platforms were released to the public, with Reddit still being one of the most popular social media platform as of 2023. Software engineers quickly started to make use of these platforms, and established Stack Exchange Network which then transformed to Stack Overflow, a Q&A website that focused on programming-related questions. Research has followed to learn more about this transition and revolution. Storey et al. well described the revolution of social media in software engineering from non-digital, digital to socially enabled (including Q&A platform) (Storey et al., 2014). Squire et al. conducted the first study to assess the quality of developer support provided by Q&A sites such as SO, in comparison with previous tools used for developer support, such as mailing list (Squire, 2015). According to Mamykina et al., some software developers started to believe that Q&A platforms such as SO had replaced web search/forums as their main source of finding answers to their programming problems (Mamykina et al., 2011a). With this trend happening, we are curious about how Q&A platforms will continue to evolve to better fit programmers’ needs. Hence we chose SO, one of the most popular Q&A sites that concentrates on programming-related questions, as our proxy for this research.
@@ 2.1.1. Stack Overflow
Created in 2008, SO quickly became the most widely used Q&A platform for software engineers in the world with about 50 million visits monthly (Nasehi et al., 2012; SEd, 2023). SO is centered around nine design decisions: voting, tags, editing, badges, karma, pre-search, search engine optimization, user interface, and critical mass (Treude et al., 2011). As such a well-developed and popular platform, SO triggered numerous researchers to conduct studies on this platform. Mamykina1 et al. examined what contributed to SO’s success, specifically engaged with the founder of this platform (Mamykina et al., 2011a). As a question-initiated system, much research focuses on analyzing the questions asked on SO, with Treude et al.’s paper in 2011 being one of the most cited works around SO (Treude et al., 2011). They summarized 7 categories of questions: how-to, discrepancy, environment, error, decision help, conceptual, and review. In addition to the high-level summary, researchers also summarized the detailed use cases of SO (ins, 2023; Antelmi et al., 2023), discussed ’how to ask the right questions’ (Ahmad et al., 2018; Yang et al., 2014), as well as analyzing the characteristics of the answers like response time (Mamykina et al., 2011a) and attributes of recognized answers (Nasehi et al., 2012). Other aspects of SO like reward system(Wang et al., 2018), novice experience (Chatterjee et al., 2020), and social norms (Cheriyan et al., 2021), were also discussed in various studies. These studies shed light on different aspects of SO’s functioning and user behavior. In addition to analyzing data and deriving insights, researchers have also developed tools to automate various aspects of this process, including classifier (Beyer et al., 2018), datasets (Baltes et al., 2018), and code search and recommendation tool (Zagalsky et al., 2012). These tools contribute to the efficiency and comprehensiveness of SO research efforts.
@@ 2.2. Generative AI for Coding
Transitioning from human-powered Q&A platforms, there is a growing trend of people increasingly embracing AI-powered tools to aid in their programming processes. According to SO’s sentiment report in 2023, 70% of the developers are already using or plan to use AI tools in their development process (sen, 2023). This trend has been significantly influenced by the advancements in LLM, particularly publicly available tools. One noteworthy example is Github Copilot, which was announced on June 29, 2021. Copilot, usually considered as an LLM-based code generation tool (Vaithilingam et al., 2022), offers various assistive features for programmers, including the conversion of code comments into executable code and auto-completion for code segments, repetitive sections, entire methods, and functions (cop, 2023).
Although it is a relatively new topic, researchers have already begun exploring programmers’ experiences with LLM-powered code generation tools. For instance, Vaithilingam et al. conducted a within-subjects user study with 24 participants to understand how programmers use and perceive Copilot (Vaithilingam et al., 2022). While these code-generation tools offer helpful assistance, they cannot fulfill all the needs typically met by traditional question-based solutions. Alongside these tools, there has also been a rising adoption of GenAI-powered chatbots for coding, a method that closely parallels traditional Q&A platforms. For instance, Rose et al. created the Programmer’s Assistant, a conversational prototype system, to explore conversational interactions rooted in code. Beyond small-scale research endeavors, the advent of ChatGPT spurred a larger trend of embracing AI chatbots for coding purposes. Launched on November 30, 2022, ChatGPT has made a substantial social impact. Users quickly began sharing their experiences and discussions on social media platforms, providing researchers with an opportunity to investigate user experiences.
Given the novelty of this topic, researchers have just begun planning studies, but there were already many interesting works that we could learn from. One primary area of interest is testing the accuracy of ChatGPT-generated answers. Kabir et al. conducted a study using SO posts as input and assessed ChatGPT’s accuracy in generating answers and found a 52% inaccuracies (Kabir et al., 2023). Despite inaccuracies, researchers still found users’ expressing willingness to use these GenAI tools (Lau and Guo, 2023).
@@
Beyond these tensions between platforms, recent research has also begun to question ChatGPT’s effectiveness in aiding coding tasks by comparing it with other methods. For example, Choudhuri et al. conducted an empirical study and discovered that ChatGPT did not enhance participants’ productivity or self-efficacy compared to traditional non-GenAI online resources (Choudhuri et al., 2024). Instead, it significantly heightened their frustration levels. While their findings highlight significant issues with GenAI coding assistance, these results are constrained by the small sample size of their lab-based experiment. We hold a more optimistic perspective and aim to observe long-term adoption of these tools in real-world settings to integrate the strengths of both methodologies. A similar view stemmed from Cheng et al., who discussed how online communities shape developers’ trust in AI tools and explored ways to leverage community features to foster appropriate user trust in AI (Cheng et al., 2022). Rather than creating opposition, our aim is to explore opportunities for collaboration and improvement based on the unique affordances of each tool. To facilitate this discussion, we will compare ChatGPT with SO and propose strategies and solutions. We anticipate that analyzing Reddit posts will uncover long-term solutions adopted by programmers in real-world settings, providing valuable insights into how we might advance our approach.
@@ 3.2. Data Filtering
During the data exploration phase, we observed an overlap in the posts on the two identified themes. For example, many posts in r/ChatgptCoding also discussed SO. Recognizing this pattern, we combined all the post data from different channels together as our final dataset. We then proceeded with initial data filtering such as eliminating duplicate posts. Subsequently, to enhance the precision of our dataset, manual filtering of irrelevant content was undertaken. This process involved researchers reviewing post content (selftext) and titles to ensure alignment. Ultimately, we reduced our dataset from over 4000 posts to 1,758 posts. The datasets are publicly available  [^1] https://doi.org/10.6084/m9.figshare.24128880.v1.
@@ 4.2. Analysis
We generated a wordcloud [^2] https://github.com/amueller/word_cloud to visually represent the narratives across all the posts [Fig. 2]. We also performed sentiment analysis on the posts using one of the state-of-the-art unsupervised sentiment prediction tools VADER [^3] https://github.com/cjhutto/vaderSentiment5 (Hutto and Gilbert, 2014). The sentiment analysis of the posts reveals a moderately positive tone throughout the dataset, with an average compound score of 0.22. Among these posts, 37.91% are categorized as positive, 53.70% as neutral, and 10.13% as negative. This distribution suggests that while the majority of the discussions are neutral or mildly positive, a significant portion still reflects varied sentiments, underlining the complexity of user experiences.
@@ 5. Qualitative Results
In order to answer RQ1: In what ways are individuals employing GenAI chatbots to enhance their coding experience?, we would like to discuss three themes: Strengths, Use Cases and Barriers. The comprehensive codebooks are available in the tables presented in the Appendix.
@@ 5.1. Strengths
Many individuals have reported highly successful experiences using ChatGPT to enhance their coding endeavors. In terms of the overall user experience, ChatGPT’s adoption of natural language interactions contributes to an exceptional barrier to entry. One user expressed this by stating, ”You will literally just describe the tool you need and you will get it.” Notably, even individuals who had previously given up on their programming pursuits expressed optimism about ChatGPT, believing it would help them ”see it through this time.” This gentle learning curve underscores ChatGPT’s potential in learning scenarios, particularly for novice users.
@@ 5.3. Barriers
One of the most prominent issues with ChatGPT’s answers is related to reliability. Specifically in coding scenarios, there have been instances where ChatGPT generated different fictitious codes each time it was asked the exact same question, undermining users’ trust in the system. To the surprise of many programmers, ChatGPT often fabricated non-existent libraries, commands, or citations. These red herrings, or false leads, can be particularly detrimental as they place a significant burden on users to verify each piece of information, potentially eroding trust in the entire solution. Users have also noted that even when ChatGPT provides incorrect or ambiguous answers, it does so with a ”calm and assured confidence,” highlighting a disparity between ChatGPT’s responses and users’ expectations, especially when users begin to treat it as a real assistant and apply human-like judgment. This unreliability is exacerbated by the fact that ChatGPT operates as a standalone program. As one user put it, ”ChatGPT only helps if you can judge whether its answer is correct, unlike an answer with 50 upvotes on SO, which you _know_ is correct.”
@@
Unsurprisingly, users have encountered limitations in terms of resource access, including constrained training sets, the absence of live resources after 2021, the inability to run ChatGPT locally, and the lack of integrated development environment (IDE) plugins. Additionally, there have been instances where ChatGPT refused to generate code, especially when specific trigger prompts were used. Users have also expressed the need for an ’adult’ version of ChatGPT with fewer restrictions, as many of the current limitations are related to ethical concerns. Other access constraints include rate limitations, regional restrictions, financial implications, downtime, and system overload. Additionally, some users expressed concerns about copyright issues related to using ChatGPT’s generated code, particularly in professional settings.
@@
We noted that many Reddit posts were based on personal experiences, but one consensus emerged: ChatGPT was perceived as significantly more user-friendly than SO. SO has a long reputation for being toxic. ”The fact is, SO has a serious problem with bullying, one that SO itself has been trying really hard to mitigate, without success so far.”, from one Reddit post. In contrast, when using Chatgpt, users did not need to worry about asking ”stupid” or repeated questions.
@@
Use Cases Personal assistant + Learning potential 7 categories (Treude et al., 2011)
@@ 6.1. Long-Term Adoption of GenAI in Coding: Reddit Posts Versus Lab Experiments
Our results complement experimental studies of coding experiences with ChatGPT and Copilot (Choudhuri et al., 2024; Vaithilingam et al., 2022). Prior work has also noted that Gen AI assistants excel at providing a starting point, are easy for beginners to use, and save programmers’ time. Perhaps the most interesting findings, however, are around differing results and conclusions. Choudhuri et al. reported negative outcomes from using ChatGPT for coding assistance, noting that it neither enhanced productivity nor self-efficacy and instead increased frustration (Choudhuri et al., 2024). However, in our analysis of Reddit posts, sentiment analysis revealed moderate positivity and the qualitative feedback from users presented a more mixed picture with both strengths and barriers.
Long-term users might have had more time to adapt and find effective ways to utilize the tool, developing specific prompting strategies like “explain certain code examples as if it were explaining to a 10 year old” mentioned in one of the posts. These prior studies focused on task completion, which, while important, does not capture the long-term sentiments about these tools, which we were able to identify in the posts. In our findings, individuals who have used ChatGPT for coding assistance in real-world scenarios greatly appreciated the respectful and patient environment it fostered, contrasting with their experiences on platforms like SO. Access and copyright issues, often overlooked in lab studies, are incredibly crucial and significantly impact long-term usage, especially in professional settings.
Programmers have more time to explore different uses of the tools in the wild, not restricted to task completion as dictated by specific study designs. Many Reddit posts indicate that programmers use ChatGPT for coding-adjacent activities, such as generating documentation, unit tests, and managing edge cases. Other use cases, such as learning, could only be captured in long-term usage with ChatGPT, where users viewed it as a coding assistant and pair programming partner, and even used it to generate quizzes. Similarly, other emergent use cases, such as employing it as a code editor or for simulating operational systems, can be detected using our methodology.